OpenStudy (anonymous):

How can I solve for x in this equation: ln(x) + ln(x-1) = 1

OpenStudy (anonymous):

\[\ln x \left( x-1 \right)=1\]

geerky42 (geerky42):

For left side of a equation, you can apply this property:\[\log(a)+\log(b) = \log(ab)\]

OpenStudy (anonymous):

I know that it first simplifies to \[\ln (x(x-1))=1\] and then to... \[x(x-1)=e\]

OpenStudy (anonymous):

solve the quadratic

geerky42 (geerky42):

Arrange, then you can apply quadratic formula.

geerky42 (geerky42):

damn you so fast @surjithayer

OpenStudy (anonymous):

Of course! Thank you both!! :D haha